# Pulse shape simulation for the reduced charge collection layer in $p$-type high-purity germanium detectors

### v1

@@ 3.3 Pulse shape validation
The numerical simulation generates pulse shapes directly. We can also generate semi-analytical pulse shapes by convolving the analytically derived distribution of hole arrival time at the $p$-$n$ boundary with a numerically simulated pulse of a single hole originating exactly at the $p$-$n$ boundary, as demonstrated in Figure 9. Figure 10 compares the numerical and semi-analytical pulse shapes for events generated at various depths in the RCC layer. Again, we see very good consistency between the simulation and analytical approaches.
@@ 4.1 Experimental setup and detector fields
We conducted a measurement on a $p$-type Broad Energy Germanium detector (BEGe) with an uncollimated 133Ba source centered at the top of the detector, see Figure 11. The $n^{+}$ electrode of the detector was heavily doped with lithium impurities via thermal diffusion. The detector parameters are listed in Table 1 dai_pulse_2024 ; dai_modeling_2023 . Figure 11 shows the r-z cross-section of the crystal. It also shows the surface sampling points that we use in the following as starting points for several charge-drift simulations. The sampling positions include points from the bottom, bottom corner, side, top corner, and top, which are used to display the results along the surface in several plots below. The figure also shows the center-to-surface angle $\theta$ that we define as
@@
Table 1: Parameters of the BEGe detector (hole and electron lifetimes are assumed to be identical in this work).
@@ 4.2 Drift paths and pulse shapes
Figure 15: The pulse shapes of events originated at the $p$-$n$ boundary (in the sensitive region) at the sampling positions which are annotated in Figure 12.
Figure 16: The pulse shapes of events originated at a depth of 0.5 mm from the surface (in the RCC layer) at the sampling positions which are annotated in Figure 12.
@@ 4.4 CCE-corrected simulated spectrum
Minor discrepancies observed for some data points (mainly ~ 70 keV), may arise from the oversimplification of the currently employed constant-lifetime trapping model, insufficient knowledge of the detector parameters, neglect of the position dependence of CCE, or neglect of the self-repulsion effect, listed in descending order of likelihood. The actual lifetime likely varies with depth in the RCC layer due to the rapid change of impurity density in the RCC layer, see Figure 2. A depth-dependent-lifetime trapping model may improve simulation precision significantly and will be the subject of future work.
@@ 5 Summary and outlook
Although this paper focuses on $p$-type HPGe detectors, the method and code are also applicable to $p$-type multi-contact HPGe detectors. Moreover, with adjustments to the RCC layer impurity profile, the method and code can also be extended to $n$-type HPGe detectors and other types of semiconductors (e.g., Silicon, CdZnTe).

### v2

@@ 3.3 Pulse shape validation
The numerical simulation generates pulse shapes directly. We can also generate semi-analytical pulse shapes by convolving the analytically derived distribution of hole arrival time at the $p$-$n$ boundary with a numerically simulated pulse of a single hole originating exactly at the $p$-$n$ boundary bjorn_lehnert_search_2016 , as demonstrated in Figure 9. Figure 10 compares the numerical and semi-analytical pulse shapes for events generated at various depths in the RCC layer. Again, we see very good consistency between the simulation and analytical approaches.
@@ 4.1 Experimental setup and detector fields
We conducted a measurement on a $p$-type Broad Energy Germanium detector (BEGe) with an uncollimated 133Ba source centered at the top of the detector, see Figure 11. The $n^{+}$ electrode of the detector was heavily doped with lithium impurities via thermal diffusion. The detector parameters are listed in Table 1 dai_pulse_2024 ; dai_modeling_2023 . The acceptor impurity density constant was derived by the experimental calibration. The hole and electron lifetimes in the RCC layer are assumed to be identical in the experimental study.
Figure 12 shows the r-z cross-section of the crystal. It also shows the surface sampling points that we use in the following as starting points for several charge-drift simulations. The sampling positions include points from the bottom, bottom corner, side, top corner, and top, which are used to display the results along the surface in several plots below. The figure also shows the center-to-surface angle $\theta$ that we define as
@@
Table 1: Parameters of the BEGe detector used in the simulation.
@@ 4.2 Drift paths and pulse shapes
Figure 15: The simulated pulse shapes of events originated at the $p$-$n$ boundary (in the sensitive region) at the sampling positions which are annotated in Figure 12.
Figure 16: The simulated pulse shapes of events originated at a depth of 0.5 mm from the surface (in the RCC layer) at the sampling positions which are annotated in Figure 12.
@@ 4.4 CCE-corrected simulated spectrum
Minor discrepancies observed for some data points (mainly ~ 70 keV) may arise from the oversimplification of the currently employed constant-lifetime trapping model, insufficient knowledge of detector parameters, or the neglect of the position dependence of CCE and self-repulsion effect, listed in descending order of likelihood.
Notably, the actual carrier lifetime is expected to vary with depth in the RCC layer, driven by a non-uniform distribution of trapping centers. This variation may originate from the crystal growth process or, more plausibly, from external contaminants introduced during fabrication. The density profile of such contaminants would likely exhibit a depth gradient, analogous to the lithium density variation shown in Figure 2. Implementing a depth-dependent-lifetime trapping model may significantly improve the simulation accuracy.
Therefore, the fitted lifetime reported here should be interpreted as an effective value subject to systematic uncertainties related to these factors. A comprehensive study addressing all these effects, including the implementation of the depth-dependent-lifetime trapping model and the variations in detector parameters, would require a massive number of simulations.
Due to the high computational cost associated with the fine grid resolution and carrier statistics in the current framework, such a study is impractical with the current framework. To address this performance bottleneck, we are currently developing a GPU-accelerated code. An extensive investigation utilizing this accelerated code is planned for future work.
@@ 5 Summary and outlook
Although this paper focuses on $p$-type HPGe detectors, the method and code are also applicable to $p$-type multi-contact HPGe detectors. Additionally, this framework provides a solid foundation for future extensions to $n$-type HPGe and other semiconductors (e.g., Silicon, CdZnTe) through adjustments to the doping profile and charge drift model.